OpenStudy (anonymous):

Mr. Nelson lost one of his student's test papers. He knows that the other 4 students scored as follows:35,24,34,27. He also knows that the average score is 29.4. what is the score on the missing paper?

OpenStudy (anonymous):

Alright, let's assume the 5 scores of Mr. Nelson to be \[S1, S2, S3, S4, S5\] Now to find the average you would ... \[\frac{ Total-of-all-the-Scores }{ Number-of -Scores}\] or \[\frac{ S1+S2+S3+S4+S5 }{ 5 }=average\] Now, you know 4 scores. Hence, let's assume that the missing score is a variable x, like so. \[\frac{ 35+24+34+27+x }{ 5 }=29.4\] From here on, it's a simple problem of solving ze equation. Cross multiply 5 with 29.4, and up the scores you have already... \[120+x=147\] Take 120 to the other side, \[x=147-120\] Which gives us 27, and your answer. \[x=27\] I hope that solved your problem. If you need any more help dont hesitate to ask. -swas.

OpenStudy (anonymous):

Thank You! That helped

OpenStudy (anonymous):

Anytime.